Mathematics 242 – Principles of Analysis Definitions Quiz 6 – March 22, 2013

Your Name:

A) (5) Give the ε, δ form of the definition of the statement “f is continuous at c.”

B) (5) Let

f(x) = \{ \begin{matrix}0 & if x < 0 \\ 1 & if x ≥ 0. \end{matrix}

What is c = lub (\{x ∈ R | f(x) ≤ \frac{1}{2}\}). Is f(c) ≤ \frac{1}{2}?